# King’s College London, U.K

### v4

@@ Abstract
We demonstrate that Assembly Theory (AT) is reliant on a method based upon statistical compression that belongs to the LZ family of compression algorithms, family which comprises popular algorithms used in ZIP, GZIP, or JPEG. Such popular algorithms have been shown to empirically reproduce the results of AT that were reported before in successful applications to separating organic from non-organic molecules and in the context of the study of selection and evolution. In particular, assembly index is a LZ compression scheme whose statistical compressibility is bounded by entropy or other traditional LZ compression schemes, such as LZ77, LZ78, or LZW. In addition, we demonstrate that AT, its pathway complexity, the assembly index, and the assembly number define compression schemes that are subsumed into the theory of algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity. Due to AT’s current lack of logical consistency in defining causality for non-stochastic processes and lack of the empirical evidence that it outperforms other complexity measures found in the literature, we conclude that the assembly index and the assembly number do not lead to an explanation or quantification of biases in generative (physical or biological) processes, including those brought about by (abiotic or Darwinian) selection and evolution.
@@
assembly theory, assembly index, pathway complexity, assembly number, algorithmic complexity, algorithmic probability, CTM, BDM, lossless compression, LZ compression, LZ77, LZ78, LZW, Lempel–Ziv–Welch, logical depth, selection, evolution.
@@ 1 Introduction
Assembly Theory (AT) has recently garnered significant attention in some academic circles and in the scientific media. Responding to how open-ended forms can emerge from matter without a blueprint, AT purports to explain and quantify selection and evolution. The central claim of Assembly Theory advanced in [1, 2] is that objects with a high assembly index “are very unlikely to form abiotically”. While this has been contested in [3], we will demonstrate the connections of Assembly Theory and its assembly index to well-established theories that the authors claim are different and unrelated to Assembly Theory [4]. We will show that the claims advanced by its authors that Assembly Theory unifies life and biology with physics [2, 5] rely on a circular argument and the use of a popular compression algorithm, with no empirical or logical support to establish deeper connections to selection and evolution not already made before [6, 7]. We demonstrate how Assembly Theory is an approximation to algorithmic complexity (Kolmogorov-Solomonoff-Chaitin complexity) [8, 9, 10, 11], denoted by $\mathbf{K}$, based on LZ compression scheme and asymptotically bounded by (Shannon) Entropy in a fashion that is similar to previous information and graph-theoretic approaches to chemical complexity [12, 13, 14, 15]. We conclude that Assembly Theory cannot offer a different or better explanation of selection or evolution than the connections already established [6].
@@ 2.1 The assembly Index is a compression algorithm of the LZ family
Despite the authors’ assertion in the Assembly Theory (AT) paper that AT and the assembly index are unrelated to algorithmic complexity, it is evident that AT is fundamentally encompassed within the realm of algorithmic complexity [16]. The assembly index, as proposed, seeks to gauge the complexity of an object based on the number of steps in its shortest copy-counting assembly pathway [2] via a procedure equivalent to LZ compression, which in turn is a computable approximation to $\mathbf{K}$.
At its core, the assembly index shares three key elements of entropy defined by Shannon himself [17], and shares the three main elements of the LZ family of compression algorithms [18, 19] that make LZ converge to entropy at the limit: the identification of repeated blocks, the usage of a dictionary containing repetitions and substitutions, and the ability to losslessly reconstruct the original object using its minimal LZ description. The parity in number of steps for compression and decompression using the assembly index effectively reduces the definition of Assembly Theory to the length of the shortest assembly pathway, which in turn is a loose upper bound of a resource-bounded approximation to algorithmic complexity $\mathbf{K}$ [16].
In contrast, more robust approximations to $\mathbf{K}$, capable of capturing blocks and other causal content within an object, have been proposed for purposes ranging from exploring cause-and-effect chains to quantifying object memory and characterising process content [20, 21]. These more advanced measures have found application in the same contexts and domains explored by Assembly Theory, encompassing tasks like distinguishing organic from non-organic molecules [15], investigating potential connections to selection and evolution [6, 7], the detection of bio and techno signatures [22, 23], and explorations into causality [24, 25]. Importantly, when applied to the data employed as evidence by AT, the more sophisticated measures consistently outperform the assembly index (see Suplementary Information and [16]). Thus, it becomes evident that Assembly Theory and its assembly index represent a considerably constrained version of compression algorithms, and a loose upper bound of $\mathbf{K}$.
@@
Any approximation to $\mathbf{K}$ that accounts for identical repetitions, including all known lossless statistical compression algorithms, can achieve equivalent or superior results, as demonstrated in the Suplementary Information and [16]. This alignment with ‘Template Program A’ effectively highlights the association of Assembly Theory with well-established principles of compression and coding theory, thereby refuting the initial claim of its authors to present a unique methodology. Moreover, the authors’ suggestion that their index may be generalised as a universally applicable algorithm for any object (including text) [26]further underscores the disconnection of Assembly Theory and their authors’ suggestion to reinvent text compression (e.g. ZIP or GZIP) from the current state of the art in the field of statistical and non-statistical compression beyond LZW.
@@
In essence, the assembly index is fundamentally underpinned by the LZ encoding of the objects it measures. This reveals that the assembly index, and consequently AT, aligns more closely with $\mathbf{K}$ and the principles of traditional information theory (Shannon Entropy), and statistical data compression than the authors are willing to acknowledge [4].
@@ 2.2 A compression algorithm is a mechanical procedure that corresponds to a physical process
The authors of AT argue that traditional data compression algorithms are an overly abstract process unsuited for modelling the construction (or assembly) of objects [26]. This view overlooks the practical and mechanistic nature of compression algorithms, particularly those in the LZ family. Since their introduction in 1977, LZ algorithms have been effectively used in detection, identification, clustering, and classification across various fields, including biology, chemistry, and medicine  [27, 15, 28], and to approximate algorithmic complexity [10, 29].
The argument presented by the proponents of Assembly Theory regarding the uncomputability of algorithmic complexity is deeply misguided. While it is true that $\mathbf{K}$ is semi-computable, computable algorithms like LZ77/LZ78/LZW have been widely used to approximate it. These algorithms have been applied in biology and chemistry, challenging the assertion that Assembly Theory represents a unique or superior approach  [15, 16]. Furthermore, in an earlier 1976 article [30], Lempel and Ziv defined an early version of their algorithm directly as a computable method to approximate algorithmic complexity, defining what is known as LZ complexity, which speaks to the initial motivation behind their now seminal compression algorithm.
The emphasis on the purported requirement of a Turing machine in the context of algorithmic complexity is a misdirected concern. A Turing machine is a synonym of ‘algorithm’ and an ‘algorithm’ is a synonym of a Turing machine, any rule or method that is algorithmic in nature, including all the computable methods in Assembly Theory, are algorithms, and, therefore, technically, Turing machines or programs that can run on a (universal) Turing machine. These facts are only a technicality unrelated to any putative advantage of Assembly Theory over other approaches to algorithmic complexity.
On the contrary, one of the distinctive features of algorithmic information theory (AIT), algorithmic complexity being one of its indexes, is that its importance and pervasiveness in mathematics, theoretical computer science, and complexity science are, in fact, established because of the invariance of its results with respect to the model of computation, whether abstract or physically implemented. More specifically, such statements from the authors of AT [26] indicate a fundamental misunderstanding of equivalence classes closed under reductions (in this case, computability classes), which are basic and pivotal concepts in computer science, complexity theory, and mathematics in general. It is evidence in support of this point that none of the formal proofs in the Sup. Inf. section A made any mention of any Turing machine.
@@
In a published paper [4], the authors offered a proof of computability of their assembly index to distance themselves from Shannon Entropy and algorithmic (Kolmogorov) complexity. Their algorithm, categorised as Template Program A, is trivially computable and requires no proof of computability but also all other resource-bounded approximations to $\mathbf{K}$ are computable, including LZW that has been used for decades for this purpose [10]
@@ 2.3 Conflating object assembly process and causation directionality
The authors assume and present the sequential nature of their algorithm as an advantage [26]. Assembly Theory claims to be able to extract causal knowledge in the form of chains of cause and effect from its assembly pathways [26]. The authors’ central assumption is that each step of the basic Template Program A constitutes a cause-and-effect chain corresponding to how an object may have been physically assembled. This is no different from, and is indeed a restricted version of, a pushdown automaton capable of instantiating grammar compressor (which the authors rename as ‘assembly pathway’) and works exactly like an LZ algorithm.
@@
This assumption of sequential assembly of an object based on causal direction is manifestly incorrect regarding how an object subject to selection assembles. In contrast, only by finding the generative mechanism of the object—such as the underlying set of (computable) mechanisms (of which the assembly index only takes into account one instance, the program that only counts copies)–and thus explaining the object in a non-trivial fashion, can one reproduce both the causal direction from the sequence of connected steps (see Fig. 1B) as well as how the object itself may have been assembled (which is not and cannot be in a sequential fashion).
Assembly Theory and its index cannot characterise this, for example, because the class of problems or functions recognised by pushdown automata (or generated by context-free grammars) is a proper subset of the class of recursive problems.
The authors of Assembly Theory have mistakenly claimed that algorithmic complexity and Turing machines are not, or cannot be, related to causality. This is incorrect [24, 31]. They were introduced as causal artifacts for studying mechanistic means for logical operations.
We have demonstrated that the assembly index requires a finite automaton to instantiate a restricted grammar (see Sup. Inf. A), a version of a specific-purpose machine, just as any other resource-bounded approximation to $\mathbf{K}$ would need to be instantiated (e.g. calculated, even by hand). In this case, the size of their formula or the size of the implementation of their computer program is the size of the special-purpose finite automaton, which is common and of fixed size for their calculations (which allows them to be discounted from the final length, just as it is from $\mathbf{K}$ or resource-bounded approximations).
The assembly index being reduced to a grammar compression scheme demonstrates not only that the assembly process’ complexity is dependent on the contingencies regarding which assembly spaces or paths were taken in the historical past stages to build the object, but also that this recursivity and dependence on past trajectories are subsumed into the concept of compressibility. Unlike algorithmic approaches, assembly theory does not incorporate any of the elements of causality such as perturbation and counterfactual analysis, as has been done in the context of algorithmic complexity, a generalisation that compared to AT reflects the simplistic nature of AT [25].
@@ 2.4 Lack of control experiments and absence of supporting empirical data beyond current domain knowledge
A critical examination of the AT methodology reveals significant shortcomings. Firstly, proponents of AT failed to conduct basic control experiments, a foundational aspect of introducing a new scientific metric. Benchmarking against established indices, particularly in coding and compression algorithms, is crucial to validating any new metric in the domain. Previous work on AT has never included meaningful experimental comparisons of the assembly index with other existing measures but schemes and mistaken theoretical arguments [4] (Figs. 1 and 2). Yet, we have shown that other algorithms, such as RLE, Huffman coding, and those based on entropy and dictionary-based methods, reproduced equivalent or superior results compared to the results published by the AT group [16].
@@
Assembly Theory introduces a cutoff value in its index, purported to offer a unique perspective on molecular complexity, distinct from those based on algorithmic complexity. This value indicates when an object is more likely to be organic, alive, or a product of a living system. However, various indices and compression algorithms tested have yielded equivalent cutoff values (see [16]). In fact, such an Assembly Index threshold replicates previous results on molecular separation that employ algorithmic complexity. Thus, this overlap in results challenges the notion that Assembly Theory provides a unique tool for distinguishing between organic and non-organic matter. For instance, similar cutoff values have been derived in other studies, such as those mentioned in  [15], effectively separating organic from non-organic molecular compounds. While the authors of AT ignored decades of research in chemical complexity based on graph theory and the principles of Entropy [13, 12, 32, 14], all of which are equivalent to Assembly Theory, the depth of the purported experimental validation of AT was notably limited, especially when contrasted with more comprehensive studies. For instance, a more exhaustive and systematic approach, with proper control experiments involving over 15,000 chemical compounds [15] and employing algorithms from the LZ family (or Entropy) and others of greater sophistication (BDM), demonstrated the ability to separate organic from non-organic molecular compounds.
@@
Turing’s motivation was to explore what could be automated in the causal mechanisation of operations (see Fig. 1B), for example, stepwise by hand using paper and pencil. The shortest among all the computer programs of this type is an upper bound approximation to $\mathbf{K}$. In other words, $\mathbf{K}$ cannot be longer than the length of this diagram. The concept of pathway complexity and copy number instantiated by the assembly index, is a restricted version of statistical nature of a Turing machine or finite automaton. Moreover, a Turing machine is a synonym of an algorithm and vice versa. It is a mistake to think of a Turing machine as a physical object or an object with particular properties or features (such as a head or tape). All the algorithms in assembly theory are Turing machines and viceversa.
@@
Figure 1: A: The authors of AT have suggested that $\mathbf{K}$ would be proven to be contained in Assembly TheoryT [26]. This Venn diagram shows how Assembly Theory (AT) is connected and subsumed within algorithmic complexity or ($\mathbf{K}$) and within the group of statistical compression as proven in this paper (see Sup. Inf. A). B: Causal transition graph of a Turing machine with number 3019 (in Wolfram’s enumeration scheme [33]) with an empty initial condition found by using a computable method (e.g. CTM, [34]) to explain how the block-patterned string 111000111000 was assembled step-by-step based on the principles of $\mathbf{K}$ describing the state, memory, and output of the process as a fully causal mechanistic explanation. A Turing machine is simply a procedural algorithm and any algorithm can be represented by a Turing machine. By definition, this is a mechanistic process, and as physical as anything else, not an ‘abstract’ or ‘unrealisable’ process.

### v5

# Assembly Theory is an approximation to algorithmic complexity based on LZ compression that does not explain selection or evolution

@@ Abstract
We demonstrate that Assembly Theory (AT) is reliant on Shannon Entropy and a method based upon statistical compression that belongs to the LZ family of compression algorithms, family which comprises popular algorithms used in ZIP, GZIP, or JPEG. Such popular algorithms have been shown to empirically reproduce the results of AT that were reported before in successful applications to separating organic from non-organic molecules and in the context of the study of selection and evolution. In particular, the assembly index calculation method is an LZ compression scheme, even if used for classification or explanatory purposes. We show that the assembly index value is equivalent to the size of a minimal context-free grammar. The statistical compressibility of such a method is bounded by Shannon Entropy and other equivalent traditional LZ compression schemes, such as LZ77, LZ78, or LZW. In addition, we demonstrate that AT, and the algorithms supporting its pathway complexity, assembly index, and assembly number, define compression schemes and methods that are subsumed into the theory of algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity. Due to AT’s current lack of logical consistency in defining causality for non-stochastic processes and lack of the empirical evidence that it outperforms other complexity measures found in the literature capable of explaining the same phenomena, we conclude that the assembly index and the assembly number do not lead to an explanation or quantification of biases in generative (physical or biological) processes, including those brought about by (abiotic or Darwinian) selection and evolution that could not have been made with Shannon Entropy or had not been reported before with algorithmic complexity.
@@
assembly theory, assembly index/number, pathway complexity, algorithmic complexity, lossless compression, LZ compression, LZ77, LZ78, logical depth, selection, evolution.
@@ 1 Introduction
Assembly Theory (AT) has recently garnered significant attention in some academic circles and the scientific media. Responding to how open-ended forms can emerge from matter without a blueprint, AT purports to explain and quantify selection and evolution. The central claim of AT advanced in [1, 2] is that objects with a high assembly index “are very unlikely to form abiotically”. This has been contested in [3], whose results “demonstrate that abiotic chemical processes have the potential to form crystal structures of great complexity”, exceeding the assembly index (or MA, when applied to quantify the assembly index on molecules) threshold proposed by AT’s authors. The existence of such abiotic objects would lead AT’s methods to be prone to false positives, corroborating the predictions in [4]. The assembly index (or MA) was shown to perform equally, if not worse in some cases, to popular compression algorithms (including those of statistical nature) [4], some of which have been applied before in [5, 6, 7].
The forthcoming sections show the connections of AT to well-established theories that AT’s authors claim are different from and unrelated to [8]. We show that the claims advanced by its authors that AT unifies life and biology with physics [2, 9] rely on a circular argument and the use of a popular compression algorithm, with no empirical or logical support to establish deeper connections to selection and evolution not already known (hierarchical modularity), already made (in connection to complexity) [6, 10, 11], or previously investigated by information- and graph-theoretic approaches to chemical and molecular complexity [12, 13, 14, 5].
The non-linear (or “tree”-like) structure of the minimum rooted assembly (sub)spaces is one of the ways to define a (compression) scheme that encodes the assembling process itself back into a linear sequence of codewords. We demonstrate that the generative/assembling process in an assembly space is strictly equivalent to a Context-Free Grammar (CFG), while the assembly index value is equivalent to the size of such a compressing grammar. This also implies that the assembly index can only be an approximation to the number of factors [15] in a traditional LZ scheme, such as LZ78 or LZW, which are statistical compression methods whose compression rates converge to that of Shannon Entropy.
In addition to the assembly index calculation method being a CFG-based compression method bounded by traditional LZ schemes, we demonstrate that the assembling process that results in the construction of an object is a LZ scheme. The compressibility achieved by this LZ scheme depends on the length of the shortest assembling paths and the “simplicity” of the minimal assembly spaces. This notion of simplicity encompasses both how the assembly space structure differs from a single-thread (or linear) space and how many more distinct assembling paths can lead to the same object. Such a LZ scheme reduces the complexity (which the assembly index aims to quantify) of the assembling process of an object to compressibility and computational resource efficiency of a compression method.
The theory and methods of AT describing a compression algorithm are, therefore, subsumed into Algorithmic Information Theory (AIT), which in turn has been applied in the same areas that AT has covered before, from organic versus non-organic compound classification to bio and technosignature detection and selection and evolution [5, 16, 6, 17]. This makes the assembly index and assembly number purported by AT effectively approximations to algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity [18, 19, 20, 21].
As proposed in [2], the assembly number is intended to measure the amount of selection and evolution necessary to produce the ensemble of (assembled) objects. That is, the assembly number aims to quantify the presence of constraints or biases in the underlying generative processes (e.g., those part of the environment in which the objects were assembled) of the ensemble, processes that set the conditions for the appearance of the assembled objects. A higher assembly number—not to be conflated with the assembly index—would mean that more “selective forces” were put into play as e.g. environmental constraints or biases in order to allow or generate a higher concentration of high-assembly-index elements. Otherwise, these high-assembly-index objects would not occur as often in the ensemble.
Consonant with the constraints and biases that the assembly number aims to quantify but that, in fact, we demonstrate it constitutes a compression method subsumed into AIT, ensembles with higher assembly numbers are more compressible and, therefore, would diverge more from those ensembles that are more statistically random with a lesser presence of constraints and biases. If one assumes that the assembly number quantifies the presence of constraints or biases in the underlying generative processes of the ensembles, then a more compressible ensemble implies that more constraints or biases played a role in generating more high-assembly-index objects more frequently than would have been the case in an environment with otherwise fewer constraints and biases (i.e., a more random or incompressible environment), hence increasing the frequency of occurrence of high-assembly-index (i.e., less compressible) objects in this environment.
Conversely, under the same assumption, in case more biotic processes are present in an ensemble, it implies a higher assembly number, and thereby it would also imply that the ensemble is more compressible. For example, this occurs in scenarios where there is a stronger presence of top-down (or downward) causation [11] behind the possibilities or paths that lead to the construction of the objects, while a less compressible ensemble would indicate a weaker presence (or absence) of top-down causation.
We, therefore, conclude that AT cannot offer a different or better explanation of selection, evolution, or top-down causation than the connections already established [6, 11] consistent with our previous position that a single scalar is unlikely to classify life or quantify selection or evolution independent of the environment and the perturbations such environment imposes on the objects and the assembly process.
@@ 2.1 The assembly index is a compression algorithm of the LZ family
Despite the authors’ assertion in the Assembly Theory (AT) paper that this theory and the assembly index are unrelated to algorithmic complexity, it is evident that AT is fundamentally encompassed within the realm of algorithmic complexity [4]. The assembly index, as proposed, seeks to gauge the complexity of an object based on the number of steps in its shortest copy-counting assembly pathway [2] via a procedure equivalent to LZ compression, which in turn is a computable approximation to algorithmic complexity, denoted by $\mathbf{K}$. See Sup. Inf. A.
At its core, the assembly index shares three key elements of entropy defined by Shannon himself [22], and shares the three main elements of the LZ family of compression algorithms [23, 24] that make LZ converge to Shannon Entropy at the limit: the identification of repeated blocks, the usage of a dictionary containing repetitions and substitutions, and the ability to losslessly reconstruct the original object using its minimal LZ description. As demonstrated in the Sup. Inf. A, the parity in number of steps for compression and decompression using the assembly index effectively reduces the definition of AT to the length of the shortest assembly pathway, which in turn is a loose upper bound of a resource-bounded approximation to algorithmic complexity $\mathbf{K}$.
In contrast, more robust approximations to $\mathbf{K}$, capable of capturing blocks and other causal content within an object, have been proposed for purposes ranging from exploring cause-and-effect chains to quantifying object memory and characterising process content [25, 26]. These more advanced measures have found application in the same contexts and domains explored by AT, encompassing tasks like distinguishing organic from non-organic molecules [5], investigating potential connections to selection and evolution [6, 10], the detection of bio and techno signatures [27, 28], and explorations into causality [29, 30]. Importantly, when applied to the data employed as evidence by AT, the more sophisticated measures consistently outperform the assembly index (see Sup. Inf. A and [4]). Thus, it becomes evident that AT and its assembly index represent a considerably constrained version of compression algorithms, and a loose upper bound of $\mathbf{K}$.
@@
Any approximation to $\mathbf{K}$ that accounts for identical repetitions, including all known lossless statistical compression algorithms, can achieve equivalent or superior results, as demonstrated in the Suplementary Information and [4]. This alignment with ‘Template Program A’ effectively highlights the association of AT with well-established principles of compression and coding theory, thereby refuting the initial claim of its authors to present a unique methodology. Moreover, the authors’ suggestion that their index may be generalised as a universally applicable algorithm for any object (including text) [31] further underscores the disconnection of AT and their authors’ suggestion to reinvent text compression (e.g. ZIP or GZIP) from the current state of the art in the field of statistical and non-statistical compression beyond LZW.
@@
In essence, the assembly index is fundamentally underpinned by the LZ encoding of the objects it measures. This reveals that the assembly index, and consequently AT, aligns more closely with the principles of traditional information theory (Shannon Entropy), and statistical data compression than the authors are willing to acknowledge [8]. In fact, both AT and the statistical compression methods that underpins it are subsumed into the theory of algorithmic complexity.
@@ 2.2 A compression algorithm is a mechanical procedure that corresponds to a physical process
The authors of AT argue that traditional data compression algorithms are an overly abstract process unsuited for modelling the construction (or assembly) of objects [31]. This view overlooks the practical and mechanistic nature of compression algorithms, particularly those in the LZ family. Since their introduction in 1977, LZ algorithms have been effectively used in detection, identification, clustering, and classification across various fields, including biology, chemistry, and medicine  [32, 5, 33], and to approximate algorithmic complexity $\mathbf{K}$ [20, 34].
The argument presented by the proponents of AT regarding the uncomputability of algorithmic complexity is deeply misguided. While it is true that $\mathbf{K}$ is semi-computable, computable algorithms like LZ77/LZ78/LZW have been widely used to approximate it. These algorithms have been applied in biology and chemistry, challenging the assertion that AT represents a unique or superior approach  [5, 4]. Furthermore, in an earlier 1976 article [35], Lempel and Ziv defined an early version of their algorithm directly as a computable method to approximate algorithmic complexity, defining what is known as LZ complexity, which speaks to the initial motivation behind their now seminal compression algorithm.
The emphasis on the purported requirement of a Turing machine in the context of algorithmic complexity is a misdirected concern. A Turing machine is a synonym of ‘algorithm’ and an ‘algorithm’ is a synonym of a Turing machine, any rule or method that is algorithmic in nature, including all the computable methods in AT, are algorithms, and, therefore, technically, Turing machines or programs that can run on a (universal) Turing machine. These facts are only a technicality unrelated to any putative advantage of AT over other approaches to algorithmic complexity.
On the contrary, one of the distinctive features of algorithmic information theory (AIT), algorithmic complexity being one of its indexes, is that its importance and pervasiveness in mathematics, theoretical computer science, and complexity science are, in fact, established because of the invariance of its results with respect to the model of computation, whether abstract or physically implemented. More specifically, such statements from the authors of AT [31] indicate a fundamental misunderstanding of equivalence classes closed under reductions (in this case, computability classes), which are basic and pivotal concepts in computer science, complexity theory, and mathematics in general. It is evidence in support of this point that none of the formal proofs in the Sup. Inf. section A made any mention of any Turing machine. In the same manner, all the mentions to Turing machines in [4] can also be equivalently replaced by any sufficiently expressive programming language running on an arbitrary computer.
@@
In a published paper [8], the authors offered a proof of computability of their assembly index to distance themselves from Shannon Entropy and algorithmic (Kolmogorov) complexity. Their algorithm, categorised as Template Program A, is trivially computable and requires no proof of computability but also all other resource-bounded approximations to $\mathbf{K}$ are computable, including LZW that has been used for decades for this purpose [20].
@@ 2.3 Conflating object assembly process and causation directionality
The authors assume and present the sequential nature of their algorithm as an advantage [31]. Assembly Theory claims to be able to extract causal knowledge in the form of chains of cause and effect from its assembly pathways [31]. The authors’ central assumption is that each step of the basic Template Program A constitutes a cause-and-effect chain corresponding to how an object may have been physically assembled. This is no different from, and is indeed a restricted version of, a pushdown automaton capable of instantiating grammar compressor (which the authors rename as ‘assembly pathway’) and works exactly like a LZ algorithm. See Sup. Inf. A.
@@
This assumption of sequential assembly of an object based on causal direction is manifestly incorrect regarding how an object subject to selection assembles. In contrast, only by finding the generative mechanism of the object—such as the underlying set of (computable) mechanisms (of which the assembly index only takes into account one instance, the program that only counts copies)—and thus explaining the object in a non-trivial fashion, can one reproduce both the causal direction from the sequence of connected steps (see Fig. 1B) as well as how the object itself may have been assembled (which is not and cannot be in a sequential fashion).
AT and its index cannot characterise this, for example, because the class of problems or functions recognised by pushdown automata (or generated by context-free grammars) is a proper subset of the class of recursive problems.
The authors of AT have mistakenly claimed that algorithmic complexity and Turing machines are not, or cannot be, related to causality. This is incorrect [29, 36]. They were introduced as causal artifacts for studying mechanistic means for logical operations.
We have demonstrated that the assembly index requires a finite automaton to instantiate a context-free grammar (see Sup. Inf. A), a version of a specific-purpose machine, just as any other resource-bounded approximation to $\mathbf{K}$ would need to be instantiated (e.g. calculated, even by hand). In this case, the size of their formula or the size of the implementation of their computer program is the size of the special-purpose finite automaton, which is common and of fixed size for their calculations (which allows them to be discounted from the final length, just as it is from $\mathbf{K}$ or resource-bounded approximations).
The assembly index calculation from the minimum rooted assembly space being reduced to a grammar compression scheme demonstrates not only that the assembly process’ complexity is dependent on the contingencies regarding which assembly spaces, or which distinct paths, were taken in the historical past stages to build the object, but also that this recursivity and dependence on past trajectories are subsumed into the concept of compressibility. However, unlike algorithmic approaches, assembly theory does not incorporate any of the elements of causality such as perturbation and counterfactual analysis, as has been done in the context of algorithmic complexity [30], a generalisation that compared to AT reflects the simplistic nature of AT.
@@ 2.4 Lack of control experiments and absence of supporting empirical data beyond current domain knowledge
A critical examination of the AT methodology reveals significant shortcomings. Firstly, proponents of AT failed to conduct basic control experiments, a foundational aspect of introducing a new scientific metric. Benchmarking against established indices, particularly in coding and compression algorithms, is crucial to validating any new metric in the domain. Previous work on AT has never included meaningful experimental comparisons of the assembly index with other existing measures but schemes and mistaken theoretical arguments [8] (Figs. 1 and 2). Yet, we have shown that other algorithms, such as RLE, Huffman coding (one of the first, or the first dictionary-based universal compression algorithm), and other compression algorithms based on dictionary-based methods Shannon Entropy, reproduced equivalent or superior results compared to the results published by the AT [4].
@@
AT also introduces a cutoff value in its index, purported to offer a unique perspective on molecular complexity, distinct from those based on algorithmic complexity. This value indicates when an object is more likely to be organic, alive, or a product of a living system. However, various indices and compression algorithms tested have yielded equivalent cutoff values (see [4]). In fact, such an Assembly Index threshold replicates previous results on molecular separation that employ algorithmic complexity. Thus, this overlap in results challenges the notion that AT provides a unique tool for distinguishing between organic and non-organic matter. For instance, similar cutoff values have been derived in other studies, such as those mentioned in [5], effectively separating organic from non-organic molecular compounds. While the authors of AT ignored decades of research in chemical complexity based on graph theory and the principles of Entropy [13, 12, 37, 14], all of which are equivalent to AT, the depth of the purported experimental validation of AT was notably limited, especially when contrasted with more comprehensive studies. For instance, a more exhaustive and systematic approach, with proper control experiments involving over 15,000 chemical compounds [5] and employing algorithms from the LZ family (or Entropy) and others of greater sophistication (BDM), demonstrated the ability to separate organic from non-organic molecular compounds.
@@
Turing’s motivation was to explore what could be automated in the causal mechanisation of operations in Fig. 1B, for example, stepwise by hand using paper and pencil. The shortest among all the computer programs of this type is an upper bound approximation to $\mathbf{K}$. In other words, $\mathbf{K}$ cannot be longer than the length of this diagram. The concept of pathway complexity and the algorithm for copy number instantiated by the assembly index, is a restricted version of a Turing machine or finite automaton. Moreover, a Turing machine is a synonym of an algorithm and vice versa. It is a mistake to think of a Turing machine as a physical object or an object with particular properties or features (such as a head or a tape). All the algorithms in assembly theory are Turing machines and viceversa.
@@
Figure 1: A: The authors of AT have suggested that $\mathbf{K}$ would be proven to be contained in AT [31]. This Venn diagram shows how AT is connected and subsumed within algorithmic complexity or ($\mathbf{K}$) and within the group of statistical compression as proven in this paper (see Sup. Inf. A). B: Causal transition graph of a Turing machine with number 3019 (in Wolfram’s enumeration scheme [38]) with an empty initial condition found by using a computable method (e.g. CTM, [39]) to explain how the block-patterned string 111000111000 was assembled step-by-step based on the principles of $\mathbf{K}$ describing the state, memory, and output of the process as a fully causal mechanistic explanation. A Turing machine is simply a procedural algorithm and any algorithm can be represented by a Turing machine. By definition, this is a mechanistic process, and as physical as anything else, not an ‘abstract’ or ‘unrealisable’ process.